OpenStudy (anonymous):

I need help in this math problem. Purvi and her friends tipped their waiter a total of $12 for dinners that cost $9 , $8, $11 , $12 and each. Use a percent proportion to solve for the percent tip they left the waiter.

OpenStudy (anonymous):

Anyone plz.

OpenStudy (anonymous):

I really need help. I got the answer, but i am not sure

OpenStudy (anonymous):

it needs to be a percent

OpenStudy (anonymous):

they spent a total of $40 and \[\12\div 40 = .3=30\%\]

OpenStudy (anonymous):

like a percent proportion

OpenStudy (anonymous):

i got 12/ 40 = x/100

OpenStudy (anonymous):

i think it is 30 %

OpenStudy (anonymous):

thx for the help

OpenStudy (radar):

Step 1. Find the total price of the meal: $(9+8+11+12) = $40.00 12= x*40 x=12/40=.3 The proportion is 12/40 as x/100 12:40 :: x:100 solve the proportion 40x=1200 x=30%

OpenStudy (anonymous):

thx

OpenStudy (radar):

you're welcome and good luck with your studies.

OpenStudy (radar):

They were generous tippers.